f(x) = x + 2 and asked to evaluate for f(-2) and f(0)
Answer by jim_thompson5910(35256)

f(x) = x + 2

f(-2) = -2 + 2

f(-2) = 0
=======================================================
f(x) = x + 2

f(0) = 0 + 2

f(0) = 2